# He awa whiria: the tidal streams of interstellar objects

### v1

@@ 1 Introduction
In this work, we consider two additional, qualitatively different, effects that differentiate the dynamics of the Galactic population of ISOs from stars, expanding the scope and sophistication of the Ōtautahi-Oxford model. First, when ISOs are unbound from their parent star system, they may find themselves in the birth cluster of their star. Such ISOs will then populate the cluster potential and, eventually, escape from it (Levison et al., 2010; Hands et al., 2019). Regardless of this potential “preprocessing” by the birth cluster, nearly all will eventually orbit in the Galactic potential. Second, any material orbiting in the Galaxy with a finite spread in either initial position or velocity will generically produce tidal tails or streams. This phenomenon is well-studied in the context of stellar streams produced by star clusters and dwarf galaxies disrupting in the Milky Way’s potential, motivated by the discovery of stellar streams in large surveys (e.g. Odenkirchen et al., 2001; Rockosi et al., 2002). Observed stellar streams are excellent tracers of the Galactic potential because of their low internal velocity dispersions (Bonaca & Price-Whelan, 2024), and they may also retain information about their interactions with dense substructures (Bonaca et al., 2019). The cumulative Galactic interstellar object population is often described with two broad assumptions to first order: that it is well-mixed, and that it is of uniform density. However, ISOs will also form streams within their Galaxy. This was observed in simulations by Portegies Zwart (2021) in one realization of their ISO populations originating from the 200 nearest stars to Earth. Radiants of small interstellar meteors could form in this way (Taylor et al., 1996; Baggaley, 2000), though lack confirmed observations (Musci et al., 2012; Froncisz et al., 2020; Hajduková et al., 2024).
@@ 2 Stream Simulations
Since we are not dealing with stellar streams, but rather ISO streams, it is worth explaining our placement of ISOs at L1 and L2 in more depth. In our model, the progenitor of an ISO stream may be a single star, or a star cluster. In the latter case, the ISOs will have been ejected from their individual parent star system, but early enough and at a low enough velocity that they remain bound to the star cluster. The ISOs will then be subject to similar perturbations that allow some stars in the cluster to escape, though the process may not be perfectly analogous given the dramatically lower masses of the ISOs. Nonetheless we expect the placement of ISOs at L1 and L2 to be an excellent approximation in this particular case. In the case of ISOs ejected from a single star, however, we need to consider two sub-cases: ISOs ejected from single encounters with a planet, or ISOs eroded from an Oort cloud (discussed more in § 4.1). Erosion from an Oort cloud, in analogy to the cluster case, is likely to occur preferentially at the saddle points of the potential, i.e. we again expect L1 and L2 to be a reasonable location to place the ISOs. ISOs ejected from the inner parts of the star system via strong encounters with a planet, however, will be ejected essentially from the progenitor rather than the Lagrange points. In this case the typical velocity of ejection will be far larger than the Keplerian velocity at the Lagrange points, and so it will be unimportant that we choose that the particle was initialized at a Lagrange point instead of the progenitor, assuming that the ejections are isotropic [^4] Depending on the details of the system, these ejections by planets may occur preferentially in the ecliptic plane of the progenitor system, which would qualitatively change the results of this section (Portegies Zwart, 2021). However, due to the subsequent influence of heating (§ 4.2), we expect that any early anisotropy in the ejection geometry will be washed out within about a Gyr.. We are therefore in a situation that the only time the particle’s initial location matters is when they escape at the Lagrange points, so it is reasonable to initialize all particles at the Lagrange points.
@@
Given a set of test particle positions and velocities at a variety of ejection times between $t_{0}$ and $t_{f}$, it is straightforward (though cumbersome) to integrate their motion forward in the Galactic potential. To avoid running these integrations explicitly, we instead simply “write down” the orbital solution following the formalism of Lynden-Bell (2015), essentially a higher-order epicyclic approximation. We use our LBparticles package666github.com/lbparticles/lbparticles, which implements an extended version of the Lynden-Bell (2015) formalism to include vertical motions under the influence of a disk with an arbitrary radial density profile, and corrects several minor errors in the cosine expansions relating time or azimuthal angle to corresponding internal angles (analogous to the mean and true anomaly for Keplerian orbits) (Forbes et al in prep). We adopt the axisymmetric time-independent MW2014 potential (Bovy, 2015) throughout this work.
With the test particle orbits now accessible via LBparticles, we can evaluate the position and velocity of all test particles in the stream at any time $t$ at minimal marginal cost. To understand the stream’s properties based on the particles’ positions and velocities at a moment in time, it is useful to construct an estimator of the stream’s phase space density $f(\vec{x},\vec{v})$. The streams generally have nontrivial six-dimensional phase space structure – they are highly anisotropic, with over- and under-dense regions, and low sinuosity. The stream develops over-dense regions due to the scale imprinted by the individual epicyclic motions of particles (Küpper et al., 2008). Moreover, dynamically cold streams will have thin substructures or “feathering” (Amorisco, 2015). In the presence of such structures, an isotropic kernel density estimator (KDE) with a kernel that does not vary from point to point may be insufficiently accurate. We employ a new adaptive KDE for this task, as detailed in Appendix A. Hyperparameters of the adaptive KDE are fixed via cross-validation, and a convergence test is shown in Appendix B.
@@
Figure 2: The structure of an interstellar object stream 10 Gyr after its formation. For a progenitor star on a Sun-like orbit, $10^{4}$ particles are generated at the star’s Lagrange points with a velocity dispersion of $3.1\ \mathrm{km}\ \mathrm{s}^{-1}$, and evolved for 10 Gyr in a Milky Way-like potential, where they wrap the Galaxy multiple times. The top left panel shows the estimated density of the stream in $r-\phi$ coordinates centered on the progenitor (green star). The sequence of “scaffolding” orbits $\vec{\mathcal{S}}_{P}$ is shown as a red line, with ticks every kpc indicating the direction of $\hat{e}_{1}$. The lower left panel shows the corresponding $r-z$ view of the stream’s density (grey) and the scaffolding orbits (red). On the right, two example locations along the stream connected by black lines the $r-\phi$ diagram display the density of a cross section of the stream (grey), centered on $\vec{\mathcal{S}}_{P}$ (red cross). For four points in each of these cross sections, the velocity structure of the stream is shown in adjacent plots (linear particle histogram in colourmap). Finally, the lower-right panel show the density of the stream per unit length along the stream $dM/d\ell$, with colored lines connecting various $\ell$ coordinates to their positions in the $r-\phi$ diagram.
@@
Figure 4: Summary of ISO stream morphologic properties. The top four panels compare the width, height, length, and peak density of a set of simulated streams to rules of thumb for each of these quantities. In the bottom row, the velocity dispersion along the 3 different axes of the stream is compared to the initial velocity dispersion $\sigma_{0}$. The bottom right panel shows the shape parameters of the density distribution. In each panel, 5 points are shown for each stream, corresponding to 5 randomly-selected times between 0 and 10 Gyr. For the most part we expect streams to follow the 1:1 line shown in these panels, and for the most part they do, over 3+ orders of magnitude. The exception is $\sigma_{L}$, the velocity dispersion along the stream direction, which is typically far less than $\sigma_{0}$, since a particle’s position in the stream is primarily determined by its velocity in this direction. Note that for plotting clarity we have included a small random scatter of 0.07 dex on the x-coordinates of points where the x-axis is $\sigma_{0}$ times a constant, since many points have identical values.
@@
We simulate a variety of streams to understand how their properties (e.g. length, width, density distribution) scale with input properties (e.g. $\sigma_{0}$, progenitor orbit, and progenitor mass). For most of our sample of simulated streams, the progenitor of each stream is a star on a Sun-like orbit, with each stream simulating a different $\sigma_{0}$. Two streams with cluster progenitors are simulated. Additionally, we explore the effect of orbital eccentricity by simulating two streams with progenitor orbits closer to circular: the thin-disk-like stream (labeled “Thin”) has a velocity difference from circular $0.31$ that of the Sun’s deviation from circular, and the stream labeled “Circ” is just on a circular orbit. Finally, the stream labeled “1 Gyr” extends the time for high rate of supply of the ISOs, setting $t_{f}-t_{0}=1\ \mathrm{Gyr}$, whereas all other streams have the short supply time of $t_{f}-t_{0}=10\ \mathrm{Myr}$.
@@ 3 Population Model
Figure 5: Interstellar object streams across the Galaxy. To visualize the wide range of streams that may intersect the Sun at any given moment, we show a population of 160 progenitors with 120 ISO tracer particles per progenitor shown in blue. The progenitors are drawn from a uniform distribution in radius. Three streams are highlighted for demonstration in other colors (purple, red, and maroon). The Sun’s orbit is shown in orange, and the streams here have evolved for 198 Myr.
The population of ISOs observed in the Solar System will originate from a variety of streams, each of which will have its own velocity dispersion, lifetime, heating history, and progenitor (cf. Figure 5). Given our understanding of the fundamental properties of streams, we now aim to predict how streams will affect the population of ISOs encountering our Solar System. At the most basic level, we would like to know how many streams contribute to the populations of ISOs visiting the Solar System. If this number is very large, each ISO is likely to come from a different stream, whereas if a few streams dominate, we would see ISO “siblings,” that is, multiple ISOs from the same stream. In the most extreme case where multiple ISOs come from a dynamically-cold stream, they may appear as “twins”: ISOs with very similar incoming velocities and radiants [^9] As far as our models are concerned, any binary ISO would be considered here as a single ISO for its dynamics in the Galaxy and encounter rate with the Sun; its components would not count as siblings in the sense we use throughout this paper.. In order to compare the relative properties of each population model in this section, we will normalize the total rate of ISOs entering a sphere of $q_{\mathrm{max}}=5$ au around the Sun to $10\ \mathrm{yr}^{-1}$. This is a broad-brush approximation to the estimates of the background number density of ISOs and the ability to find ISOs of a decade-long survey such as LSST. A more detailed accounting would depend on ISO orbits, size distributions, albedos, and observing strategy (Dorsey et al. in prep).
@@
To account for these effects, we construct a Monte Carlo model to draw ISOs from the Galactic population. We assume that each progenitor generates a number of ISOs proportional to the mass of the progenitor, and that these ISOs form a stream. For now, we neglect the complication that each stream may be composed of multiple components. Rather, we take each stream to be isothermal, characterized by a single age [^11] Note that this is slightly different than our setup in the simulations of individual streams, where we initialized particles at a range of times — for the population models, we assume the ISOs are released in a single burst for simplicity. $\tau$ that we take to be equal to the time between the birth of the star or cluster and the present day, a velocity dispersion $\sigma$, a progenitor velocity $\vec{v}_{\mathrm{prog}}$, a progenitor mass $M_{\mathrm{prog}}$, and three numbers characterizing the stream’s generalized gamma density distribution (Table 1). Draws from the Monte Carlo distribution proceed as follows. First, a set of progenitor properties is drawn. The mass of the progenitor is drawn from a Kroupa (2001) initial mass function (IMF) for the single star models, or for the cluster cases, a power law distribution between $M_{\mathrm{min}}$ and $M_{\mathrm{max}}$ with slope $\beta$. The typical velocity of the stream in the local standard of rest is taken to be a 3D isotropic Gaussian with 1D velocity dispersion $\sigma_{\mathrm{prog}}$; this distribution can be made more realistic in the future following Hopkins et al. (2024). The age of the stream is drawn uniformly from 0-12 Gyr, reflecting the roughly constant star formation history expected for the Milky Way (e.g. Schönrich & Binney, 2009). Note that the progenitor star or cluster does not need to still be extant for the ISO stream to persist, so it is the star formation history, not the age distribution of current stars, that matters (Lintott et al., 2022; Hopkins et al., 2023).

### v2

@@ 1 Introduction
In this work, we consider two additional, qualitatively different, effects that differentiate the dynamics of the Galactic population of ISOs from stars, expanding the scope and sophistication of the Ōtautahi-Oxford model. First, when ISOs are unbound from their parent star system, they may find themselves in the birth cluster of their star. Such ISOs will then populate the cluster potential and, eventually, escape from it (Levison et al., 2010; Hands et al., 2019). Regardless of this potential “preprocessing” by the birth cluster, nearly all will eventually orbit in the Galactic potential. Second, any material orbiting in the Galaxy with a finite spread in either initial position or velocity will generically produce tidal tails or streams. This phenomenon is well-studied in the context of stellar streams produced by star clusters and dwarf galaxies disrupting in the Milky Way’s potential, motivated by the discovery of stellar streams in large surveys (e.g. Odenkirchen et al., 2001; Rockosi et al., 2002). Observed stellar streams are excellent tracers of the Galactic potential because of their low internal velocity dispersions (Bonaca & Price-Whelan, 2024), and they may also retain information about their interactions with dense substructures (Bonaca et al., 2019). The cumulative Galactic interstellar object population is often described with two broad assumptions to first order: that it is well-mixed, and that it is of uniform density. However, ISOs will also form streams within their Galaxy. This was observed in simulations by Portegies Zwart (2021) in one realization of their ISO populations originating from the 200 nearest stars to Earth. Radiants of small interstellar meteors could form in this way (Taylor et al. (1996); Baggaley (2000); Gregg & Wiegert (2025)), though lack confirmed observations (Musci et al., 2012; Froncisz et al., 2020; Hajduková et al., 2024).
@@ 2 Stream Simulations
Since we are not dealing with stellar streams, but rather ISO streams, it is worth explaining our placement of ISOs at L1 and L2 in more depth. In our model, the progenitor of an ISO stream may be a single star, or a star cluster. In the latter case, the ISOs will have been ejected from their individual parent star system, but early enough and at a low enough velocity that they remain bound to the star cluster. The ISOs will then be subject to similar perturbations that allow some stars in the cluster to escape (Flammini Dotti et al., 2025), though the process may not be perfectly analogous given the dramatically lower masses of the ISOs. Nonetheless we expect the placement of ISOs at L1 and L2 to be an excellent approximation in this particular case. In the case of ISOs ejected from a single star, however, we need to consider two sub-cases: ISOs ejected from single encounters with a planet, or ISOs eroded from an Oort cloud (discussed more in § 4.1). Erosion from an Oort cloud, in analogy to the cluster case, is likely to occur preferentially at the saddle points of the potential, i.e. we again expect L1 and L2 to be a reasonable location to place the ISOs. ISOs ejected from the inner parts of the star system via strong encounters with a planet, however, will be ejected essentially from the progenitor rather than the Lagrange points. In this case the typical velocity of ejection will be far larger than the Keplerian velocity at the Lagrange points, and so it will be unimportant that we choose that the particle was initialized at a Lagrange point instead of the progenitor, assuming that the ejections are isotropic [^4] Depending on the details of the system, these ejections by planets may occur preferentially in the ecliptic plane of the progenitor system, which would qualitatively change the results of this section (Portegies Zwart, 2021). However, due to the subsequent influence of heating (§ 4.2), we expect that any early anisotropy in the ejection geometry will be washed out within about a Gyr.. We are therefore in a situation that the only time the particle’s initial location matters is when they escape at the Lagrange points, so it is reasonable to initialize all particles at the Lagrange points.
@@
Given a set of test particle positions and velocities at a variety of ejection times between $t_{0}$ and $t_{f}$, it is straightforward (though cumbersome) to integrate their motion forward in the Galactic potential. To avoid running these integrations explicitly, we instead simply “write down” the orbital solution following the formalism of Lynden-Bell (2015), essentially a higher-order epicyclic approximation. We use our LBparticles package666github.com/lbparticles/lbparticles, which implements an extended version of the Lynden-Bell (2015) formalism to include vertical motions under the influence of a disk with an arbitrary radial density profile, and corrects several minor errors in the cosine expansions relating time or azimuthal angle to corresponding internal angles (analogous to the mean and true anomaly for Keplerian orbits) (Forbes et al in prep). While this method includes vertical oscillation of the particles with a frequency that depends on galactocentric distance, it does not include the effect discussed in Dehnen & Hasanuddin (2018), namely that the vertical oscillation frequency depends on the amplitude of the oscillation leading to tidal ribbons rather than streams. This effect is moderated provided that the vertical oscillations of particles remain well within a disk scaleheight. We adopt the axisymmetric time-independent MW2014 potential (Bovy, 2015) throughout this work.
With the test particle orbits now accessible via LBparticles, we can evaluate the position and velocity of all test particles in the stream at any time $t$ at minimal marginal cost. To understand the stream’s properties based on the particles’ positions and velocities at a moment in time, it is useful to construct an estimator of the stream’s phase space density $f(\vec{x},\vec{v})$. The streams generally have nontrivial six-dimensional phase space structure – they are highly anisotropic, with over- and under-dense regions, and low sinuosity. The stream develops over-dense regions due to the scale imprinted by the individual epicyclic motions of particles (Küpper et al., 2008). Moreover, dynamically cold streams will have thin substructures or “feathering” (Amorisco, 2015). More realistic streams will have additional structures imprinted by resonances with Galaxy-scale non-axisymmetric features like the bar (Pearson et al. (2017)), and/or compact massive structures (Bonaca et al., 2019). In the presence of structure in the stream’s phase space density, even in axisymmetric time-independent potentials, an isotropic kernel density estimator (KDE) with a kernel that does not vary from point to point may be insufficiently accurate. We employ a new adaptive KDE for this task, as detailed in Appendix A. Hyperparameters of the adaptive KDE are fixed via cross-validation, and a convergence test is shown in Appendix B.
@@
Figure 2: The structure of an interstellar object stream 10 Gyr after its formation. For a progenitor star on a Sun-like orbit, $10^{4}$ particles are generated at the star’s Lagrange points with a velocity dispersion of $3.1\ \mathrm{km}\ \mathrm{s}^{-1}$, and evolved for 10 Gyr in a Milky Way-like potential, where they wrap the Galaxy multiple times. The top left panel shows the estimated density of the stream in $r-\phi$ coordinates centered on the progenitor (green star). The sequence of “scaffolding” orbits $\vec{\mathcal{S}}_{P}$ is shown as a red line, with ticks every kpc indicating the direction of $\hat{e}_{1}$. The lower left panel shows the corresponding $r-z$ view of the stream’s density (grey) and the scaffolding orbits (red). On the right, two example locations along the stream connected by black lines the $r-\phi$ diagram display the density of a cross section of the stream (grey), centered on $\vec{\mathcal{S}}_{P}$ (red cross). For four points in each of these cross sections, the velocity structure of the stream is shown in adjacent plots (linear particle histogram in colourmap). Finally, the lower-right panel shows the density of the stream per unit length along the stream $dM/d\ell$, with thin colored regions showing the $\ell$ range covered by the 5 rectangles in the $r$-$\phi$ and $r$-$z$ diagrams.
@@
Figure 4: Summary of ISO stream morphologic properties. The top four panels compare the width, height, length, and peak density of a set of simulated streams to rules of thumb for each of these quantities. In the bottom row, the velocity dispersion along the 3 different axes of the stream is compared to the initial velocity dispersion $\sigma_{0}$. The bottom right panel shows the shape parameters of the density distribution. In each panel, 5 points are shown for each stream, corresponding to 5 randomly-selected times between 0 and 10 Gyr. For the most part we expect streams to follow the 1:1 line shown in these panels, and for the most part they do, over 3+ orders of magnitude. The exception is $\sigma_{L}$, the velocity dispersion along the stream direction, which is typically far less than $\sigma_{0}$, since a particle’s position in the stream is primarily determined by its velocity in this direction. Note that for plotting clarity we have included a small random scatter of 0.07 dex on the x-coordinates of points where the x-axis is $\sigma_{0}$ times a constant, since many points have identical values. In the last panel comparing the shape parameters of the generalized gamma distribution, we include dashed grey lines at $y=x/2$, $y=x$, and $y=2x$.
@@
We simulate a variety of streams to understand how their properties (e.g. length, width, density distribution) scale with input properties (e.g. $\sigma_{0}$, progenitor orbit, and progenitor mass). For most of our sample of simulated streams, the progenitor of each stream is a star on a Sun-like orbit, with each stream simulating a different $\sigma_{0}$. Two streams with cluster progenitors are simulated. Additionally, we explore the effect of orbital eccentricity by simulating two streams with progenitor orbits closer to circular: the thin-disk-like stream (labeled “Thin Disk”) has a velocity difference from circular $0.31$ that of the Sun’s deviation from circular, and the stream labeled “Circular” is just on a circular orbit. A stream with triple the Sun’s peculiar velocity is also simulated and labelled “Thick Disk”. Finally, the streams labeled “1 Gyr” and “4.4 Gyr” extend the time for high rate of supply of the ISOs, setting $t_{f}-t_{0}=1\ \mathrm{Gyr}$ or $4.4$ Gyr respectively, whereas all other streams have the short supply time of $t_{f}-t_{0}=10\ \mathrm{Myr}$.
@@ 3 Population Model
Figure 5: Interstellar object streams across the Galaxy. To visualize the wide range of streams that may intersect the Sun at any given moment, we show a small representative population of 160 progenitors (the black points) with 120 ISO tracer particles per progenitor shown in blue. This is a top-down view of the Galaxy, and the circles show circles of constant galactocentric radius in intervals of 2 kpc. The progenitors are drawn from a uniform distribution in radius. Three streams are highlighted for demonstration in other colors (purple, red, and maroon). The Sun’s orbit is shown in orange, and the streams here have evolved for 40 Myr (left) and 198 Myr (right). The orbits of the particles are highly idealized, since the potential (MW2014 from Bovy (2015)) is axisymmetric and does not evolve with time. The most dramatic effect missing from this visualization is likely that particles within the region affected by the bar ($r\lesssim 4$ kpc; Bland-Hawthorn & Gerhard (2016)) will be redistributed by resonances with the bar’s orbit (e.g., Athanassoula (2013)).
The population of ISOs observed in the Solar System will originate from a variety of streams, each of which will have its own velocity dispersion, lifetime, heating history, and progenitor (cf. Figure 5). Given our understanding of the fundamental properties of streams, we now aim to predict how streams will affect the population of ISOs encountering our Solar System. At the most basic level, we would like to know how many streams contribute to the populations of ISOs visiting the Solar System. If this number is very large, each ISO is likely to come from a different stream, whereas if a few streams dominate, we would see ISO “siblings,” that is, multiple ISOs from the same stream. In the most extreme case where multiple ISOs come from a dynamically-cold stream, they may appear as “twins”: ISOs with very similar incoming velocities and radiants [^9] As far as our models are concerned, any binary ISO would be considered here as a single ISO for its dynamics in the Galaxy and encounter rate with the Sun; its components would not count as siblings in the sense we use throughout this paper.. In order to compare the relative properties of each population model in this section, we will normalize the total rate of ISOs entering a sphere of $q_{\mathrm{max}}=5$ au around the Sun to $10\ \mathrm{yr}^{-1}$. This is a broad-brush approximation to the estimates of the background number density of ISOs and the ability to find ISOs of a decade-long survey such as LSST. A more detailed accounting would depend on ISO orbits, size distributions, and observing strategy (Dorsey et al. (2025)).
@@
To account for these effects, we construct a Monte Carlo model to draw ISOs from the Galactic population. We assume that each progenitor generates a number of ISOs proportional to the mass of the progenitor, and that these ISOs form a stream. For now, we neglect the complication that each stream may be composed of multiple components. Rather, we take each stream to be isothermal, characterized by a single age [^11] Note that this is slightly different than our setup in the simulations of individual streams, where we initialized particles at a range of times — for the population models, we assume the ISOs are released in a single burst for simplicity. $\tau$ that we take to be equal to the time between the birth of the star or cluster and the present day, a velocity dispersion $\sigma$, a progenitor velocity $\vec{v}_{\mathrm{prog}}$, a progenitor mass $M_{\mathrm{prog}}$, and three numbers characterizing the stream’s generalized gamma density distribution (Table 1). Draws from the Monte Carlo distribution proceed as follows. First, a set of progenitor properties is drawn. The mass of the progenitor is drawn from a Kroupa (2001) initial mass function (IMF) for the single star models, or for the cluster cases, a power law distribution between $M_{\mathrm{min}}$ and $M_{\mathrm{max}}$ with slope $\beta$ so that the number of clusters per unit mass $dN/dM\propto M^{\beta}$ for $M_{\mathrm{min}}<M<M_{\mathrm{max}}$. The typical velocity of the stream in the local standard of rest is taken to be a 3D isotropic Gaussian with 1D velocity dispersion $\sigma_{\mathrm{prog}}$; this distribution can be made more realistic in the future following Hopkins et al. (2024). It is important to note that the distribution of velocities in the Solar neighborhood is likely to be affected by resonances with the bar (e.g. Moreno et al. (2021); Dillamore et al. (2024)). Our adopted density structure for each stream (see §2), which may also be affected by the bar (e.g. Pearson et al. (2017)), may therefore need to be altered for certain parts of the local velocity distribution. We expect that this will be a second-order effect (that is, adopting a velocity distribution that more closely matches the Gaia data following Hopkins et al. (2024) will be more important), but given that the origins of the structures in velocity space are not well-understood, these effects may matter more than we expect. The age of the stream is drawn uniformly from 0-12 Gyr, reflecting the roughly constant star formation history expected for the Milky Way (e.g. Schönrich & Binney, 2009). Note that the progenitor star or cluster does not need to still be extant for the ISO stream to persist, so it is the star formation history, not the age distribution of current stars, that matters (Lintott et al., 2022; Hopkins et al., 2023).